# Beyond GPT-5: Making LLMs Cheaper and Better via Performance–Efficiency Optimized Routing

### v1

@@
Abstract
Balancing performance and efficiency is a central challenge in large language model (LLM) advancement. GPT-5 addresses this with test-time routing, dynamically assigning queries to either an efficient or a high-capacity model during inference. In this work, we present Avengers-Pro, a test-time routing framework that ensembles LLMs of varying capacities and efficiencies, providing a unified solution for all performance-efficiency tradeoffs. The Avengers-Pro embeds and clusters incoming queries, then routes each to the most suitable model based on a performance-efficiency score. Across 6 challenging benchmarks and 8 leading models—including GPT-5-medium, Gemini-2.5-pro, and Claude-opus-4.1—Avengers-Pro achieves state-of-the-art results: by varying a performance-efficiency trade-off parameter, it can surpass the strongest single model (GPT-5-medium) by +7% in average accuracy. Moreover, it can match the average accuracy of the strongest single model at 27% lower cost, and reach $\sim$90% of that performance at 63% lower cost. Last but not least, it achieves a Pareto frontier, consistently yielding the highest accuracy for any given cost, and the lowest cost for any given accuracy, among all single models. Code is available at https://github.com/ZhangYiqun018/AvengersPro.
Figure 1: Avengers-Pro optimizes the trade-off between performance (accuracy) and efficiency (cost). (A) By varying a trade-off parameter $\alpha$, Avengers-Pro establishes a Pareto frontier. Compared to all single models, it achieves the highest accuracy for any given cost, and achieves the lowest cost for any given accuracy. (B) With comparable cost, Avengers-Pro outperforms the strongest single model GPT-5-medium by $7.1\%$. With comparable performance, Avengers-Pro achieves a $26.9\%$ cost reduction compared to GPT-5-medium.
1 Introduction
@@
“GPT-5 is a unified system with a smart, efficient model that answers most questions, a deeper reasoning model (GPT-5 thinking) for harder problems, and a real-time router that quickly decides which to use based on conversation type, complexity … ”
@@
In this work, we advance test-time routing to optimize the performance–efficiency trade-off. We build upon our earlier work Avengers zhang2025avengers —which showed that a simple routing recipe using ten models ($\sim$7B parameters each) surpass GPT-4.1 and 4.5 across 15 datasets—and introduce the Avengers-Pro. With a focus on performance-efficiency trade-off, the Avengers-Pro operates through three lightweight operations: (i) embedding: encode queries using a text embedding model, (ii) clustering: group queries by semantic similarity, and (iii) scoring: evaluate models within each cluster based on a performance-efficiency score weighted by a trade-off parameter $\alpha$. During inference, each query is embedded and mapped to its top-$p$ nearest clusters. The model with the highest performance-efficiency score aggregated over those clusters is selected to generate the response.
@@
In our experiments, the Avengers-Pro consists of 8 models from 4 families: GPT-5-chat, GPT-5-medium, Claude-4.1-opus, Claude-4-sonnet, Gemini-2.5-pro, Gemini-2.5-flash, Qwen3-235B-A22B-thinking-2507, and Qwen3-235B-A22B-2507. We evaluate the Avengers-Pro on 6 challenging benchmarks: GPQA-Diamond rein2024gpqa , Human’s Last Exam phan2025humanity , HealthBench arora2025healthbench , ARC-AGI chollet2024arc , SimpleQA wei2024measuring , LiveCodeBench jain2024livecodebench , and $\tau$2-bench barres2025tau . We find that compared to the strongest single model GPT-5-medium (average accuracy: $62.25\%$, cost: \$47.96\\$47.96), the Avengers-Pro can attain $7\%$ performance gain with a comparable cost (average accuracy: $66.66$, cost: \$47.13\\$47.13), and cut $27\%$ cost with a comparable performance (average accuracy: $62.66$, cost: \$35.05\\$35.05). By varying the trade-off parameter $\alpha$, the Avengers-Pro achieves an even more favorable balance between performance and efficiency. For example, to reach 90% of GPT-5-medium’s performance—a level comparable to Gemini-2.5-pro—the Avengers-Pro reduces cost by 63% relative to GPT-5-medium and by 81% relative to Gemini-2.5-pro. Furthermore, we observe that the Avengers-Pro achieves a Pareto frontier: for any fixed cost, it consistently delivers the highest performance among all models at that expenditure. Conversely, for any fixed performance target, it provides the lowest cost compared to other models attaining the same accuracy.
2 Routing for Performance-Efficiency Trade-off
@@ 3 Experiments
3.1 Experimental Settings
Models
@@ 3.1 Experimental Settings
Benchmarks
  • GPQA-Diamond rein2024gpqa : A graduate-level google-proof Q&A benchmark.
  • Human’s Last Exam (HLE) phan2025humanity : A frontier multi-modal benchmark of closed-ended academic questions. In this study, we use the text-only setting without custom patches, tool use, or retrieval during evaluation. For efficiency and reproducibility, we use the first 500 questions from the released pool and report accuracy under the official evaluation protocol.
  • ARC-AGI chollet2024arc : A benchmark focused on fluid intelligence, testing the ability to reason and solve novel problems. We use the first 200 questions from the released pool and report accuracy under the official evaluation protocol.
  • SimpleQA wei2024measuring : A factuality benchmark for short, fact-seeking questions. We use the official implementation with the default configuration and report accuracy under the official scoring. We evaluate on a subset of 500 examples uniformly sampled from the released dataset.
  • LiveCodeBench jain2024livecodebench : A dynamic, contamination-controlled coding benchmark that continuously ingests newly released problems. We evaluate on the latest public release (v6) using the official implementation and evaluation harness with the default configuration, without custom patches or post-processing.
  • $\tau^{2}$-bench barres2025tau : A controlled testbed for agents that must reason effectively and guide user actions.
For all benchmarks, we use the official repositories/implementations with their recommended parameter settings.
Implementation Details
We use k-means clustering with $k=60$ clusters. Each query is encoded by the Qwen3-embedding-8B zhang2025qwen3embeddingadvancingtext into a 4,096-dimensional semantic vector. Following common practice in routing routerdc ; zhuang2024embedllmlearningcompactrepresentations ; zhang2025avengers , we randomly split the data: 70% is used to fit the clustering model and estimate per-cluster statistics, and the remaining 30% is reserved for routing and evaluation. At inference time, we compute the embedding of the incoming query and retrieve the top-$p$ nearest clusters ($p=4$) in the embedding space. For each model $i$, we then sum its cluster-wise cost–capability scores $q_{j}^{i}$ over these three clusters and select the model with the highest total to generate the response.
Table 1: Model cost information (OpenRouter).
@@
Table 2: Benchmark information.
@@
ARC-AGI-v1 chollet2024arc pass@1 200
GPQA-Diamond rein2024gpqa pass@1 198
HLE phan2025humanity pass@1 500
LiveCodeBench-v6 jain2024livecodebench pass@1 1,055
$\tau^{2}$-bench barres2025tau pass@1 150
SimpleQA wei2024measuring pass@1 500
@@
3.2 Results and Analysis
Avengers-Pro outperforms top single models.
Avengers-Pro achieves a superior performance-efficiency trade-off.
Table 3: The performance and efficiency of Avengers-Pro vs. single models. Note that GPT-5-chat has no score on the $\tau^{2}$-bench benchmark because this model does not support tool calling. Bold indicates the best performance of a given benchmark. With $\alpha=0.1$, Avengers-Pro, surpasses GPT-5-medium in average accuracy with a $7\%$ performance gain. With $\alpha=0.53$, it matches GPT-5-medium’s average accuracy, while cutting the cost by $27\%$. With $\alpha=0.39$, it reaches 90% of GPT-5-medium’s performance at a $63\%$ lower cost.
@@ 3.2 Results and Analysis
Effects of the trade-off parameter
As shown in Figure 2, we gradually increase the trade-off parameter $\alpha$, placing more weight on performance over efficiency. As $\alpha$ increases, the average accuracy increases rapidly for small $\alpha$ and then plateaus near $\alpha\!\approx\!0.6$. On the other hand, as $\alpha$ increases, cost remains low until about $\alpha\!\approx\!0.4$ before rising sharply. These trends reveal two elbows (around $0.4$ and $0.6$) that offer favorable trade-offs.
Figure 2: Effects of the trade-off parameter $\alpha$ on the performance and efficiency. A greater value of $\alpha$ prioritizes performance over efficiency. The increase in performance is usually accompanied the increase in cost.
Figure 3: Proportion of model usage, given various trade-off parameters $\alpha$. When $\alpha$ is low, Avengers-Pro tend to route queries to Qwen3 and Qwen3-thinking. With a greater value of $\alpha$, Avengers-Pro favors GPT5-medium and Qwen3-thinking.
Proportion of model usage
As shown in Figure 3, when $\alpha$ is low, Avengers-Pro tends to favor the Qwen3 and Qwen3-thinking model, routing a great proportion of queries to these two models with a low unit price. As $\alpha$ increases, the usage of GPT-5-medium rises rapidly; concurrently, the usage of Gemini-2.5-pro and Claude-opus-4.1, which excel at complex reasoning but have a higher unit price, also increases.
4 Conclusions

### v2

@@
Abstract.
Balancing performance and efficiency is a central challenge in large language model (LLM) advancement. GPT-5 addresses this with test-time routing, dynamically assigning queries to either an efficient or a high-capacity model during inference. In this work, we present Avengers-Pro, a test-time routing framework that ensembles LLMs of varying capacities and efficiencies, providing a unified solution for all performance-efficiency tradeoffs. The Avengers-Pro embeds and clusters incoming queries, then routes each to the most suitable model based on a performance-efficiency score. Across 6 challenging benchmarks and 8 leading models—including GPT-5-medium, Gemini-2.5-pro, and Claude-opus-4.1—Avengers-Pro achieves state-of-the-art results: by varying a performance-efficiency trade-off parameter, it can surpass the strongest single model (GPT-5-medium) by +7% in average accuracy. Moreover, it can match the average accuracy of the strongest single model at 27% lower cost, and reach $\sim$90% of that performance at 63% lower cost. Last but not least, it achieves a Pareto frontier, consistently yielding the highest accuracy for any given cost, and the lowest cost for any given accuracy, among all single models. Code is available at https://github.com/ZhangYiqun018/AvengersPro. **footnotetext: Equal contribution.footnotetext: Corresponding author.footnotetext: Project lead, hushuyue@pjlab.org.cn, zhangyiqun344@gmail.com
Large Language Models, Model Routing, Cost-effective, Multi-objective Optimization
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3772429.3772445conference: Make sure to enter the correct conference title from your rights confirmation email; NOV 21–24, 2025; Londonccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
Figure 1. Avengers-Pro optimizes the trade-off between performance (accuracy) and efficiency (cost). (A) By varying a trade-off parameter $\alpha$, Avengers-Pro establishes a Pareto frontier. Compared to all single models, it achieves the highest accuracy for any given cost, and achieves the lowest cost for any given accuracy. (B) With comparable cost, Avengers-Pro outperforms the strongest single model GPT-5-medium by $7.1\%$. With comparable performance, Avengers-Pro achieves a $26.9\%$ cost reduction compared to GPT-5-medium.
1. Introduction
@@
“GPT‑5 is a unified system with a smart, efficient model that answers most questions, a deeper reasoning model (GPT‑5 thinking) for harder problems, and a real‑time router that quickly decides which to use based on conversation type, complexity … ”
@@
In this work, we advance test-time routing to optimize the performance–efficiency trade-off, and introduce the Avengers-Pro. Given a set of models and a set of labeled query-answer pairs, the Avengers-Pro operates through three lightweight operations: embedding, clustering and scoring. Specifically, first, it encodes the queries from the dataset sing a text embedding model, and then clusters them based on their semantic representations. Next, to assess each model’s capabilities and efficiency, it evaluates each model on the dataset and computes a performance-efficiency score for each cluster. Weighted by a trade-off parameter $\alpha$, this score reflects both a model’s performance (measured by its accuracy on the queries within a cluster) and its efficiency (quantified by the cost incurred when answering those queries). During inference, each query is embedded and mapped to its top-$p$ nearest clusters. The model with the highest performance-efficiency score aggregated over those clusters is selected to generate the response.
@@
In our experiments, the Avengers-Pro consists of 8 models from 4 families: GPT-5-chat, GPT-5-medium, Claude-4.1-opus, Claude-4-sonnet, Gemini-2.5-pro, Gemini-2.5-flash, Qwen3-235B-A22B-thinking-2507, and Qwen3-235B-A22B-2507. We evaluated Avengers–Pro on 6 challenging benchmarks: GPQA-Diamond (Rein et al., 2024), Human’s Last Exam (Phan et al., 2025), HealthBench (Arora et al., 2025), ARC-AGI (Chollet et al., 2024), SimpleQA (Wei et al., 2024), LiveCodeBench (Jain et al., 2024), and $\tau$2-bench (Barres et al., 2025). We find that compared to the strongest single model GPT-5-medium (average accuracy: $62.25\%$, cost: $\mathdollar 47.96$), the Avengers-Pro can attain $7\%$ performance gain with a comparable cost (average accuracy: $66.66$, cost: $\mathdollar 47.13$), and cut $27\%$ cost with a comparable performance (average accuracy: $62.66$, cost: $\mathdollar 35.05$). By varying the trade-off parameter $\alpha$, the Avengers-Pro achieves an even more favorable balance between performance and efficiency. For example, to reach 90% of GPT-5-medium’s performance—a level comparable to Gemini-2.5-pro—the Avengers-Pro reduces cost by 63% relative to GPT-5-medium and by 81% relative to Gemini-2.5-pro. Furthermore, we observe that the Avengers-Pro achieves a Pareto frontier: for any fixed cost, it consistently delivers the highest performance among all models at that expenditure. Conversely, for any fixed performance target, it provides the lowest cost compared to other models attaining the same accuracy.
Note that Avengers-Pro is not the first study to explore test-time routing. As we will discuss in detail later, there is a growing line of research that leverages router-based methods to harness collective intelligence from multiple models (Chen et al., 2024a; Zhuang et al., 2024a; Zhang et al., 2025c). While most works in this area focus primarily on improving overall performance (Chen et al., 2024a; Zhuang et al., 2024a), a few recent studies have started to investigate the trade-off between performance and efficiency by routing among smaller and larger models (Jitkrittum et al., 2025; Hu et al., 2024). Compared to these prior studies, the Avengers-Pro stands out for its simplicity and effectiveness. Most previous approaches rely on training additional neural networks and require retraining when incorporating new models. In contrast, Avengers-Pro involves only three lightweight operations, requires no neural network training, and is straightforward to implement. Moreover, it is highly reproducible and does not depend on hand-crafted prompts. To incorporate newly available models, it only requires an incremental evaluation of the new models on the dataset. Despite its simplicity, the Avengers-Pro is remarkably effective. To our knowledge, this is the first demonstration that test-time routing can be used to surpass state-of-the-art proprietary single models in terms of both efficiency and performance.
2. Related Work
Figure 2. The Avengers Pro, a unified framework for dynamic routing, optimizes performance-efficiency trade-offs by intelligently ensembling language models.
The harnessing of collective intelligence from multiple models constitutes one of the frontiers of AI and ML research, and has recently attracted much interest (Lu et al., 2024a; Guo et al., 2024; Zhang et al., 2025a; Subramaniam et al., 2025; Wan et al., 2025; Zheng et al., 2025). Existing approaches in this area generally fall into three paradigms: router-based, mixture-based, and merging-based methods. This work is most closely aligned with the router-based paradigm.
The main goal of most router-based methods is to enhance the overall performance of a set of smaller models through query routing; a neural network-based router is often trained to select, for each incoming query, the model most capable of handling it (Chen et al., 2023; Shnitzer et al., 2023; Ong et al., 2024; Feng et al., 2025; Shen et al., 2024; Huang et al., 2025). LLM-Blender (Jiang et al., 2023) utilizes pairwise comparisons to select the top-$k$ models for each query and then fuses their outputs to enhance performance. ZOOTER (Lu et al., 2024b) proposes reward-guided query routing, employing tag-based label enhancement to improve training stability. RouterDC (Chen et al., 2024b), which employs dual contrastive learning to enhance routing accuracy, and EmbedLLM (Zhuang et al., 2024a) leverages learned compact model embeddings along with query embeddings to predict routing correctness. Additionally, Model-SAT (Zhang et al., 2025d) generates capability instructions from model aptitude outcomes and employs text-aligned embeddings to guide a lightweight LLM in selecting optimal candidate models. More ecently, few recent studies have also explored routing that account for the trade-off between performance and computational cost. Routellm (Ong et al., 2024) trains a binary classifier using preference data to dynamically route queries to either a stronger or a weaker LLM during inference. GraphRouter (Feng et al., 2025) constructs a heterogeneous graph comprising task, query, and LLM nodes, and predicts the performance-cost score via an edge prediction mechanism. RouterBench (Hu et al., 2024) introduces a suite of benchmarks. Our work is most closely related to (Zhang et al., 2025c) and (Jitkrittum et al., 2025), which both employ clustering in their routers. While the former focuses only on performance, the latter considers model efficiency to a be universal cost across different tasks. The Avengers-Pro is the first study to show that test-time routing can be used to surpass state-of-the-art proprietary single models in terms of both efficiency and performance.
Driven primarily by efficiency, another related line of research explore single-model hybrid reasoning methods that trim latency and token cost by adjusting the amount or stages of “thinking” within one model (e.g., Qwen3 “thinking” mode  [^2] https://huggingface.co/Qwen/Qwen3-235B-A22B and DeepSeek-V3.1  [^3] https://huggingface.co/deepseek-ai/DeepSeek-V3.1), or learn to allocate a reasoning budget via policies(Han et al., 2025) or test-time guidance(Jiang et al., 2025). Representative approaches include adaptive “think/non-think” modes, hybrid-reasoning training pipelines, and token-budget or inference-aware policies. While such methods can cut redundant CoT tokens or speed up simple cases, they cannot realize cross-model complementarity by design (all computation remains within a single backbone), and their effectiveness is often modest (Han et al., 2025; Kang et al., 2025; Alomrani et al., 2025).
3. Routing for Performance-Efficiency Trade-off
@@ 4. Experiments
4.1. Experimental Settings
4.1.1. Models
@@
  • Google: Gemini-2.5-flash (Comanici et al., 2025), Gemini-2.5-Pro (Comanici et al., 2025).
  • Anthropic: Claude-4.1-opus (Claude, 2025a), Claude-4-sonnet (Claude, 2025b).
  • OpenAI: GPT-5-chat (OpenAI, 2025) [^4] GPT-5-chat: points to the GPT-5 snapshot currently used in ChatGPT; OpenAI recommends gpt-5 for most API usage, while gpt-5-chat exposes the latest improvements for chat use cases. Do not support function/tool call now. , GPT-5-medium (OpenAI, 2025) [^5] GPT-5-medium: GPT-5 is OpenAI’s flagship model for coding, reasoning, and agentic tasks across domains. GPT-5-medium denotes GPT-5 with reasoning_effort=medium..
  • Qwen: Qwen3-235B-A22B-2507 (or Qwen3) (Yang et al., 2025), Qwen3-235B-A22B-thinking-2507 (or Qwen3-thinking) (Yang et al., 2025).
@@
Table 1. Model cost information (OpenRouter).
@@
4.1.2. Benchmarks
Table 2. Benchmark information.
@@ 4.1.2. Benchmarks
ARC-AGI-v1 (Chollet et al., 2024) pass@1 200
GPQA-Diamond (Rein et al., 2024) pass@1 198
HLE (Phan et al., 2025) pass@1 500
LiveCodeBench-v6 (Jain et al., 2024) pass@1 1,055
$\tau^{2}$-bench (Barres et al., 2025) pass@1 150
SimpleQA (Wei et al., 2024) pass@1 500
@@
  • GPQA-Diamond (Rein et al., 2024): A graduate-level google-proof Q&A benchmark.
  • Human’s Last Exam (HLE) (Phan et al., 2025): A frontier multi-modal benchmark of closed-ended academic questions. In this study, we use the text-only setting without custom patches, tool use, or retrieval during evaluation. For efficiency and reproducibility, we use the first 500 questions from the released pool and report accuracy under the official evaluation protocol.
  • ARC-AGI (Chollet et al., 2024): A benchmark focused on fluid intelligence, testing the ability to reason and solve novel problems. We use the first 200 questions from the released pool and report accuracy under the official evaluation protocol.
  • SimpleQA (Wei et al., 2024): A factuality benchmark for short, fact-seeking questions. We use the official implementation with the default configuration and report accuracy under the official scoring. We evaluate on a subset of 500 examples uniformly sampled from the released dataset.
  • LiveCodeBench (Jain et al., 2024): A dynamic, contamination-controlled coding benchmark that continuously ingests newly released problems. We evaluate on the latest public release (v6) using the official implementation and evaluation harness with the default configuration, without custom patches or post-processing.
  • $\tau^{2}$-bench (Barres et al., 2025): A controlled testbed for agents that must reason effectively and guide user actions, we randomly sampled 50 examples from each of the three categories.
For all benchmarks, we use the official implementations with their recommended prompts and hyperparameters; for context length, we set each model to its documented maximum context window and apply no additional truncation, chunking, or sliding-window processing.
Table 3. The performance and efficiency of Avengers-Pro vs. single models. Note that GPT-5-chat has no score on the $\tau^{2}$-bench benchmark because this model does not support tool calling. Bold indicates the best performance of a given benchmark, and underline indicates the second-best. With $\alpha=0.1$, Avengers-Pro, surpasses GPT-5-medium in average accuracy with a $7\%$ performance gain. With $\alpha=0.53$, it matches GPT-5-medium’s average accuracy, while cutting the cost by $27\%$. With $\alpha=0.39$, it reaches 90% of GPT-5-medium’s performance at a $63\%$ lower cost.
@@
4.1.3. Implementation Details
We use k-means clustering with $k=60$ clusters. Each query is encoded by the Qwen3-embedding-8B (Zhang et al., 2025b) into a 4,096-dimensional semantic vector. Following common practice in routing (Chen et al., 2024a; Zhuang et al., 2024b; Zhang et al., 2025c), we randomly split the data: 70% is used to fit the clustering model and estimate per-cluster statistics, and the remaining 30% is reserved for routing and evaluation. At inference time, we compute the embedding of the incoming query and retrieve the top-$p$ nearest clusters ($p=4$) in the embedding space. For each model $i$, we then sum its cluster-wise cost–capability scores $q_{j}^{i}$ over these three clusters and select the model with the highest total to generate the response.
4.2. Results and Analysis
4.2.1. Avengers-Pro outperforms top single models.
Figure 3. Effects of the trade-off parameter $\alpha$ on the performance and efficiency. A greater value of $\alpha$ prioritizes performance over efficiency. The increase in performance is usually accompanied the increase in cost.
4.2.2. Avengers-Pro achieves a superior performance-efficiency trade-off.
Figure 4. Proportion of model usage, given various trade-off parameters $\alpha$. When $\alpha$ is low, Avengers-Pro tend to route queries to Qwen3 and Qwen3-thinking. With a greater value of $\alpha$, Avengers-Pro favors GPT5-medium and Qwen3-thinking.
4.2.3. Effects of the trade-off parameter
As shown in Figure 3, we gradually increase the trade-off parameter $\alpha$, placing more weight on performance. Accuracy (left) climbs quickly for small $\alpha$ and then saturates around $\alpha\!\approx\!0.6$; the shaded 95% CI also narrows as $\alpha$ grows, indicating more stable outcomes once the router consistently calls stronger models. Cost (right) stays near its minimum up to $\alpha\!\approx\!0.4$, then rises steeply before tapering off. The gray dashed baselines mark GPT-5-medium (62.25% accuracy; \$47.96 cost): our curve exceeds the accuracy baseline already at moderate $\alpha$ while remaining below the cost baseline across a broad range before very high $\alpha$. Taken together, the curves reveal two elbows—around $\alpha\!\approx\!0.4$ and $\alpha\!\approx\!0.6$—that correspond to regime changes in routing: (i) efficiency-first ($\alpha\!\lesssim\!0.4$), where the router mostly selects cheaper models and yields low cost with moderate accuracy; (ii) a transition band ($0.4\!<\!\alpha\!<\!0.6$), where accuracy improves rapidly per marginal dollar as the router begins invoking stronger models on harder clusters; and (iii) performance-first ($\alpha\!\gtrsim\!0.6$), where accuracy is near its ceiling and additional $\alpha$ mainly buys cost increases.
4.2.4. Proportion of model usage
As shown in Figure 4, when $\alpha$ is low, Avengers-Pro tends to favor the Qwen3 and Qwen3-thinking model, routing a great proportion of queries to these two models with a low unit price. As $\alpha$ increases, the usage of GPT-5-medium rises rapidly; concurrently, the usage of Gemini-2.5-pro and Claude-opus-4.1, which excel at complex reasoning but have a higher unit price, also increases. Consistent with Figure 1, model usage correlates with proximity to Avengers-Pro’s Pareto frontier: models closer to the frontier (GPT-series, Qwen3-series) see higher utilization, whereas those farther away (claude-series, gemini-series) are selected less frequently for a given $\alpha$.
5. Conclusions